OpenStudy (anonymous):

Aimi is making Valentine’s Day cards for everyone in her class. She plans to use a whole sheet of paper for each of her 6 closest friends. She will use 1/8 of a sheet of paper for everyone else in the class. She has 34 sheets of paper. How many of the smaller cards can she make?

OpenStudy (theeric):

Hi! I'm tired, but I'll give it a shot! If this is for a class, what class is it for? Algebra?

OpenStudy (anonymous):

yes its for algebra

OpenStudy (theeric):

Alright! Are you required to express the answer as an equation, or anything special like that? Otherwise, we can work it out more... common sensically. Rather than mathspeak.

OpenStudy (anonymous):

no nothing special... just show work..

OpenStudy (theeric):

Okay, so we'll make sure we write down all of the math. Well, I see from the information in the problem that we know how many sheets of paper we have (34), and how many we're already using for "closest friends" (6). How many sheets are left? I'm sure you can, but I have to ask.. Can you answer that? If you can, we'll figure out what you did mathematically to get there. Key word: mathematically.

OpenStudy (theeric):

Once we start using math symbols to write down what we're doing, we're into the work you have to show..

OpenStudy (anonymous):

28 sheets are left

OpenStudy (theeric):

Yep, and that was achieved with "34 - 6", so you can write that down in your work. Make sure to add labels if your teacher really likes that.

OpenStudy (ganpat):

lets do it little simple.. 6 + x/8 = 34.. x/8 = 28 x = 28 * 8 = 224 small greetings...

OpenStudy (theeric):

Then we have the confusing part. Okay... I was beaten to the punch.

OpenStudy (ganpat):

@theEric : Hope you dont mind... :)

OpenStudy (theeric):

And that's the mathy way. An equation that describes the situation will always be useful if you can come by it. \[x \times (1 \div 8) = 28\] will resolve to the same answer. As long as buddhababy002 understands as much as she wants to, I'm good.

OpenStudy (anonymous):

yeah i think i understand how to do them with the equation there.. thank you both! :D

OpenStudy (theeric):

I was going to torture buddhababy002, really. Once buddhababy002 understood the problem common sensically, I was going to go into how you'd approach math by itself. Your welcome! Thanks for thanking us! When your faced with any algebra problem where they say there's a total and tell you all the parts that add up to that total, simply add up the parts and set it equal to the total.

OpenStudy (theeric):

I think that's usually a good starting point at least. Good luck all.

OpenStudy (ganpat):

and Happy Valentine`s day !! :D :-P

OpenStudy (theeric):

Hahaha!! Same to you both!